An athlete takes a time of 9.82 s to cover a distance of 100 m. Calculate the average speed of he athlete.



1
Expert's answer
2021-11-17T10:37:30-0500
"v_{avg}=\dfrac{d}{t}=\dfrac{100\ m}{9.82\ s}=10.2\ \dfrac{m}{s}."